OpenStudy (anonymous):

Q:f={(0,5),(1,4),(2,7),(3,14)} and g(x)=3f(x+1)+7 a)evalute g(-1) b) evalute g (1) c) evalute g(3)

OpenStudy (anonymous):

a)22 b)28 c)is notdefined

OpenStudy (anonymous):

i noe but how did u solve it can u show me tht

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

a) g(-1)=3f(-1+1)+7=3f(0)+7 from the definition of f we see that f(0)=5 so g(-1)=3f(0)+7=3(5)+7=22

OpenStudy (anonymous):

g(x) = 3f(-1 + 1) + 7 g(x) = 3f(0) + 7 find f(0) from the Q:F (0,5) g(x) = 3*5 + 7 g(x) = 15 + 7 g(x) = 22 g(x) = 3f(1 + 1) + 7 g(x) = 3f(2) + 7 find f(2) = (2,7) g(x) = 3*7 + 7 g(x) = 21 + 7 g(x) = 28 g(x) = 3f(3 + 1) + 7 g(x) = 3f(4) + 7 find f(4) = not listed don't know what to do

OpenStudy (anonymous):

f is explicitly defined as the given set of ordered pairs, therefore there is nothing to "do" in part (c) , it is simply not defined

OpenStudy (anonymous):

thannks :)